OpenStudy (anonymous):

multiple choice Q1: y=2x^2+9 y=-4x+5 A. (-4,21), (-0.5, 7) B. (0.5, 3), (4,-11) C. (-4, 21), (4,61) D. (-4, 21), (0.5, 12) (This stuff is about linear system of equation)

OpenStudy (anonymous):

also if you could show work

jimthompson5910 (jim_thompson5910):

y=2x^2+9 -4x+5=2x^2+9 ... plug in y =-4x+5 0 = 2x^2 + 9 + 4x - 5 0 = 2x^2 + 4x + 4 2x^2 + 4x + 4 = 0 I'll let you finish